OpenStudy (anonymous):

Theorem: A line parallel to one side of a triangle divides the other two proportionately.

OpenStudy (anonymous):

In the figure below, segment DE is parallel to segment BC and segment EF is parallel to AB. Which statement can be proved true using the given theorem? Segment BF = 16 Segment BD = 20 Segment BD = 15 Segment BF = 32

OpenStudy (jim766):

look at the two sides that touching DE Since we that DE is parallel to BC we know that the 2 sides touching DE are proportional

OpenStudy (jim766):

a proportion is when two fractions are equal to each other, so.... 24/20 = 18/x do you know how to solve a proportion?

OpenStudy (anonymous):

nope

OpenStudy (jim766):

cross multiply 24/20 = 18/x 24x = 360

OpenStudy (anonymous):

now what?

OpenStudy (jim766):

get the x by itself.... any ideas?

OpenStudy (anonymous):

15

OpenStudy (jim766):

good... now look at the picture....what the line segment under the 18 on the left side?